OpenStudy (anonymous):

Someone please help! :) (x^2-4)/(x-8)/(x-2) (x^2+2x+1)/(x-2) over (x^2-1)/(x^2-4)

OpenStudy (anonymous):

factor your terms first

OpenStudy (anonymous):

So (x+2)(x-2)(x-2)/x-8; etc

OpenStudy (anonymous):

Wait what?? haha

OpenStudy (cwtan):

1st part isit \[\frac{(x^2-4)}{\frac{x-8}{(x-2}}\]?

OpenStudy (anonymous):

yes

OpenStudy (cwtan):

or \[\frac{(\frac{x^2-4}{x-8})}{x-2}\] it has their difference

OpenStudy (anonymous):

the first one

OpenStudy (cwtan):

so @treeman5823 has gave u the answer..... Let say 0.5=1/2 \[\frac{2}{\frac{1}{2}}\] 2/0.5=4 right? That how we move we move the 2 in the 1/2 to the numerator that's it

OpenStudy (anonymous):

ohh ok